A spiral spring extends by 5cm under a load of 60N. When the load is replaced by a steel block, the new extension is 7cm. What is the weight of the steel block
1
Expert's answer
2019-05-27T11:06:20-0400

The Hooke's law states

"F=kx"

So


"\frac{F_2}{F_1}=\frac{x_2}{x_1}"

"F_2=F_1\frac{x_2}{x_1}"

"=60\:\rm{N}\times\frac{7\:\rm{cm}}{5\:\rm{cm}}=84\:\rm{N}"